Question 1. Nimbus, Inc., makes brooms and then sells them doorto-door. Here is the relationship between the number of workers and Nimbus’s output in a given day: Workers Output Marginal Product Total Cost Average Total Cost Marginal Cost Workers Output Marginal Product Total Cost Average Total Cost Marginal Cost 0 0 1 20 2 50 3 90 4 120 5 140 6 150 7 155 a) Fill in the column of marginal products. What pattern do you see? How might you explain it? b) A worker costs $100 a day, and the firm has fixed costs of $200. Use this information to fill in the column for total cost. c) Fill in the column for average total cost. (Recall that ATC=TC/Q.) What pattern do you see? d) Now fill in the column for marginal cost. (Recall that MC=ΔTC/ΔQ.) What pattern do you see? e) Compare the column for marginal product and the column for marginal cost. Explain the relationship. f) Compare the column for average total cost and the column for marginal cost. Explain the relationship.


1
Expert's answer
2021-12-27T08:59:07-0500

Solution:


a)


"Marginal\space product= \frac{Change\space in\space output}{Change\space in\space worker}"





The marginal product initially increases till 3 units of worker and after this marginal product starts to decrease.

The reason for this pattern is due to return to a factor and law of variable proportion.


b)


Worker cost= 100

Fixed cost= 200

Total cost= 200+100(number of worker)





c)


"ATC = \frac{ TC}{output}"





ATC initially decreases and after 5 units of workers, it starts to rise. Here ATC is U-shaped


d)


"Marginal \;cost=\frac{Change\;in\;TC}{Change\;in\;Q}"





MC initially decreases and after 90 units of output, it starts to rise. Here MC is U-shaped


a)




Diminishing marginal returns hold that increase in variable input units holding a certain amount of fixed inputs, the output initially grows at a faster rate, then at a steady rate, but ultimately, it will grow at a declining rate.


 Nimbus experienced diminishing marginal product starting from adding the fourth worker.